a man of mass 60 kg start falling from building of height 80 m with a bag of 2kg in his hand after falling through a distance 20 m he throw a bad horizontally with respect to him so that he fall in a pond 2m away from vertical lines of fall calculate the velocity given to the bag and horizontal distance of bag from  the vertical line of fall where it lands :askiitians

Grade:

1 Answers

Shobhit Varshney IIT Roorkee
9 years ago
Hi,

Using, equation of motion, s = ut + 0.5at2
TIme to fall first 20m height = 2s
Time to fall the rest of the height = 2s

Horizontal velocity of man = 2m/2s = 1 m/s
Using conservation of momentum, 60 * 1 = 2 * vb
vb= 30 m/s = velocity given to bag
Horizontal distance of bag from vertical line of fall = 30 * 2 = 60 m

thanks